1/1

N. N. 82,  N. N. 82,  N. N. 82,
Venu paco  May there be peac  Hevenu shalom
translated by Amri Wandel  translated by N. N. 84

Bonvenu paco sur vin Bonvenu paco sur vin Bonvenu paco sur vin Bonvenu paco sur vin Paco sur vin May there be peace in the world… peace, peace, peace, peace all over the world.

Hevenu shalom alechem, Hevenu shalom alechem, Hevenu shalom alechem, Hevenu shalom alechem, Shalom alechem

Translation of the hebrea poem “Hevenu shalom ” by N. N. 82 into Esperanto by Amri Wandel.

Arg-998-2018 (2014-04-27 13:37:20)

Translation of the hebrea poem “Hevenu shalom ” by N. N. 82 into English by N. N. 84.

Arg-998-2017 (2014-04-27 13:29:55)

Author of this hebrea poem is N. N. 82.

Arg-998-2015 (2014-04-27 13:43:04)

Pri la teksto vidu la retejon http: // www. israelmagazin. de/ israel-juedisch/ hevenu-shalom-alechem . La kanto estas aŭdebla en http: // www. youtube. com/ watch? v= CQN_7TlikFM .

http://www.poezio.net/version?poem-id=998&version-id=2018,2017,2015